OpenStudy (anonymous):

Find the distance between the given points. C(7, 1) and D(4, -6) Distance =

OpenStudy (queelius):

 The distance formula is derived from Pythagorean's theorem, x^2 + y^2 = z^2 -> z = sqrt(x^2 + y^2). Where x is the difference between the x components of the two points, and y is the difference between the y components of the two points.